OpenStudy (anonymous):

if y=x^3-x^2-7x-3 what is the maximum value of y on the interval [-3,3]

OpenStudy (anonymous):

can you find the critical points?

OpenStudy (anonymous):

@shamil98 Do you know how to do this problem?

OpenStudy (shamil98):

Yes, you would use the chain rule, right?

OpenStudy (anonymous):

Nope, @shamil98 Is learning calculus so I want to know if this is up to his level.

OpenStudy (shamil98):

\[\frac{ df }{ dx } =\frac{ df }{ dg } \times \frac{ dg }{ }\]

OpenStudy (shamil98):

dg/dx*

OpenStudy (anonymous):

@shamil98 It is just a polynomial, so you only need to use \(x^n\to nx^{n-1}\)

OpenStudy (shamil98):

that's the formula atleast, i'm still a beginner at this stuff ._.

OpenStudy (shamil98):

you would just find the derivative?

OpenStudy (anonymous):

Well, what does derivative have to do with this problem?

OpenStudy (anonymous):

We need it to find critical points: \(y'=0\)

OpenStudy (shamil98):

Oh, i haven't gone up that point yet

OpenStudy (anonymous):

Interesting, yet you knew how to find derivative. That is amazing

OpenStudy (shamil98):

I'm not going by a class or anything, I was researching calculus and found derivatives and anti derivatives i just thought they were interesting lol

OpenStudy (e.mccormick):

@shamil98 You look at any of those books I linked last night?

OpenStudy (shamil98):

I'm in pre-calculus at the moment, but I'm taking a look at calculus through a textbook and online stuff.

OpenStudy (anonymous):

@sherief Can you find derivative?

OpenStudy (shamil98):

Sorry, mick. Could you link them again? I'm currently using "Thomas' Calculus Early Transcendental , Annotated Instructor's Edition:

OpenStudy (e.mccormick):

Very good ones there for learning how to deal with problems like this one. http://www.aimath.org/textbooks/textbooklist.html This is another good resource: http://tutorial.math.lamar.edu/Classes/CalcI/CalcI.aspx

OpenStudy (shamil98):

thanks! I'll take a look at them. Calculus is a really interesting type of math , i should've taken a look at this subject earlier ..

OpenStudy (e.mccormick):

For example, in the Paul's notes there is this: http://tutorial.math.lamar.edu/Classes/CalcI/MinMaxValues.aspx Which deals directly with the critical points Wio is going over.